OpenStudy (anonymous):

y=2.5x 2y+3x=32 Solve???

OpenStudy (anonymous):

replace the \(y\) in \(2y+3x=32\) by \(2.5x\) and solve \[2\times 2.5x+3x=32\]

OpenStudy (anonymous):

better known as \[5x+3x+32\]

OpenStudy (gudden):

You messed up while typing your second comment @satellite73 it's 5x + 3x = 32

OpenStudy (anonymous):

so i did

OpenStudy (anonymous):

Thanks @gudden @satellite73 :)

OpenStudy (anonymous):

yw